OpenStudy (anonymous):

What is a solution of x² + 6x = –5?

OpenStudy (anonymous):

Okay, so first, we want the right side of the equation to equal 0. Do you know how to do that?

OpenStudy (anonymous):

x^2+6x=-5 => x^2+6x+9=-5+9=4. This is legit since 9 is a real number. Then x^2+6x+9=(x+3)^2=4. We take the root of each side and obtain x+3=2 or x+3=-2. Hence x=-1 or x=-5. Or you can get some geometric help from Descartes.

OpenStudy (anonymous):

... or you could do it that way :D

OpenStudy (anonymous):

(x+ ?)(x+ ?) = 0 (after adding 5 both sides) where ?*? = 5